OpenStudy (anonymous):

Write an equation of the line through each point. Use slope-intercept form (1,-1); parallel to y=2/5x-3

OpenStudy (texaschic101):

do you know what the slope of the line y = 2/5x - 3 ?

OpenStudy (anonymous):

Isn't the slope 2/5?

OpenStudy (texaschic101):

correct...and a parallel line will have the same slope. Now we put it in point slope form : y - y1 = m(x - x1) slope(m) = 2/5 (1,-1) x1 = 1 and y1 = -1 now sub your info into the formula and what do you get ?

OpenStudy (anonymous):

y+-1=2/5(3-1)

OpenStudy (texaschic101):

let me get you started... y - y1 = m(x - x1) y - (-1) = 2/5(x - 1) y + 1 = 2/5(x - 1) can you finish and put it in y = mx + b form

OpenStudy (anonymous):

I thought it would be -1=2/5(3-1)

OpenStudy (texaschic101):

no...... y + 1 = 2/5(x + 1) y + 1 = 2/5x + 2/5 y = 2/5x + 2/5 - 1 y = 2/5x + 2/5 - 5/5 y = 2/5x - 3/5 or y = mx + b -1 = 2/5(1) + b and solve for b -1 = 2/5 + b -1 - 2/5 = b -3/5 = b so your equation would be : y = 2/5x - 3/5 you can do it either way

OpenStudy (anonymous):

Oh, okay. I kind of get it. Is that all?

OpenStudy (texaschic101):

yes...your equation is y = 2/5x - 3/5